OpenStudy (anonymous):

Can someone please walk me through solving this problem? I have really been struggling with it. Which choice is equivalent to the fraction below when x is an appropriate value? Hint: Rationalize the denominator and simplify. 5 / 5 + √10x ( 5 divided by 5 plus the square root of 10x)

OpenStudy (anonymous):

is it \[\large \frac{5}{5+\sqrt{10x}}\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok then multiply top and bottom by \[5-\sqrt{10x}\] and this works because \[(a+b)(a-b)=a^2-b^2\] so \[5+\sqrt{10x})(5-\sqrt{10x})=25-10x\]

OpenStudy (anonymous):

the denominator will be \(25-10x\) and the numerator will be \[5(5+\sqrt{10x})\]

OpenStudy (anonymous):

Yes I have done that already and my final answer was 5 - √10x / 5 - √10x. I got this by trying to simplify so I divided 5 by 5 and got just 5 - √10x then i did the same to the bottom but it was 25 divided by 5 and i got 5 - √10x.

OpenStudy (anonymous):

i.e. you will get \[\frac{5(5+\sqrt{10x})}{25-10x}\] now you can factor a \(5\) out of the bottom and cancel

OpenStudy (anonymous):

ok i think i have confused you lets go slow

OpenStudy (anonymous):

Okay sounds good! :)

OpenStudy (anonymous):

What first?

OpenStudy (anonymous):

\[\large \frac{5}{5+\sqrt{10x}}\] is the expression

OpenStudy (anonymous):

the conjugate of \(5+\sqrt{10x}\) is \(5-\sqrt{10x}\) so we write \[\large \frac{5}{5+\sqrt{10x}}\times \frac{5-\sqrt{10x}}{5-\sqrt{10x}}\]

OpenStudy (anonymous):

leave the numerator in factored form i.e. do NOT multiply out out

OpenStudy (anonymous):

you get \[\large \frac{5(5-\sqrt{10x})}{25-10x}\] is it clear how i got the denominator?

OpenStudy (anonymous):

Yes very clear! I can do it that far just get confused when I try and simplify.

OpenStudy (anonymous):

ok that was the hard part, now it the easy part factor the denominator \[\large \frac{5(5-\sqrt{10x})}{5(5-2x)}\]

OpenStudy (anonymous):

How did you get 2x?

OpenStudy (anonymous):

\[\large \frac{\cancel5(5-\sqrt{10x})}{\cancel5(5-2x)}=\frac{5-\sqrt{10x}}{5-2x}\]

OpenStudy (anonymous):

oh how did i get the \(2x\) ?

OpenStudy (anonymous):

the denominator was \(25-10x\) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

each term has a common factor of \(5\) right?

OpenStudy (anonymous):

Yes so did you just factor out 5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[25-10x=5(5-2x)\] is all

OpenStudy (anonymous):

which as math teachers would say "cancels" with the \(5\) in the numerator

OpenStudy (anonymous):

Okay! Yay! That was what I was missing the entire time. So in any problem similar to this one if they can both be factored by the same number do I do that? Or was it just this one?

OpenStudy (anonymous):

\[\large \frac{5(5-\sqrt{10x})}{25-10x}=\large \frac{5(5-\sqrt{10x})}{5(5-2x)}=\large \frac{\cancel5(5-\sqrt{10x})}{\cancel5(5-2x)}=\frac{5-\sqrt{10x}}{5-2x}\]

OpenStudy (anonymous):

this was cooked up to do that in general nothing cancels

OpenStudy (anonymous):

now i have a question

OpenStudy (anonymous):

Okay shoot.

OpenStudy (anonymous):

do you really go to an Aviation Academy?

OpenStudy (anonymous):

Yes I do. Ha. Are you asking because a person going to an aviation school should know how to do this?

OpenStudy (anonymous):

no because you said in your profile you were i don't want the pilot of my 787 worrying about square roots ever just concentrate on getting there safely!

OpenStudy (anonymous):

ooooh i see it is a high school where attitude meets altitude lol

OpenStudy (anonymous):

Haha very true! Although I don't want to be a pilot. I am going into the medical field and one of my dreams is to fly for AeroMed or something similar. Which is still aviation but I have not decided if I want to fly the helicopter or be the RN in it. Thank you so much for helping me out! I really appreciate it.

OpenStudy (anonymous):

yw